OpenStudy (anonymous):

What are the steps to solve this problem? 0.000027 ÷ 0.000009

OpenStudy (anonymous):

convert both to fractions and then use the process for dividing fractions..

OpenStudy (anonymous):

all the steps

OpenStudy (anonymous):

0.000027=27/1000000 0.000009 = 9/1000000

OpenStudy (anonymous):

cant i have to explaine it to my teacher

OpenStudy (anonymous):